Journalist Report – February 22nd

Crew 205 Crew Journalist Report 22-FEB-2019 Sol 12

Author: Maria Grulich

Title: “I will survive with unreasonably sized pancakes”

Two weeks have passed. Two weeks of EVAs, laughter, intercultural conversations, puzzles and a lot of coffee and large pancake in American style.

The International Emerging Space Leaders Crew 205 finished the simulation at 1700 today.

What would you do on your last day on Mars? We actually used it to clean the hab. All Crew- Heinzelmännchen were busy vacuuming, mobbing, brushing or packing. As a reward we got mac and cheese cooked deliciously as always from Hannah. The crew journalist Maria also used the time to take the official crew pictures and finishing up the outreach project which will have its glorious premiere tonight exclusively for MDRS. The commander Natalia was busy reviewing and finalizing the mission summary. From 8 strangers who have barely met before we became a team and our friendships will survive past Mars.

The Hab is blinking and the outside is calling so at 1700 sharp the crew members went outside for a little walk around the Hab as far as the ground conditions allowed.

We’d just like to wish a final, special thanks to some of the great people and Martians that have helped us during our time here.

To Hannah, for keeping the crew incredibly well fed, reading an entire cookbook, and generally giving us all the soul and spirit to enjoy our time here to the fullest

To David and Atila for being our mysterious guardian angels, and always appearing out of the fog over the hills to fix the Hab when we couldn’t.

It was a great experience for everyone and we will always remember our two weeks on Mars.

We want to leave you with our favorite song:

I will survive….

Ad Astra!

Operations Report – February 22nd

Crew 205 Operations Report 22 – FEB – 2019

SOL: 12

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1730 (22ndFeb), turned off at 1100 (22ndFeb)

Solar— SOC % – Turned off Gen (22ndFeb) 74%, Turned on Gen (22ndFeb) 94%

Diesel Reading – 70%

Propane Reading – ~ 30%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 24%, ~ 133 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 166.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01414903 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Greenhab Report – February 22nd

Crew 205 Green Hab Report – 22-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (2 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.1°C

High: 32.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 6.5 gallons (an extra 6 gallons were used for cleaning rovers)

Water in Blue Tank: ~ 166.5 gallons

Crops watered at 0900 hours and 1800 hours:

1) 10.30 AM 21.3° C / 25 %
2) 12.00 PM 19.4° C / 21 %
3) 15:00 PM 32.0° C / 11 %
4) 18:00 PM 23.4° C / 54 %
Changes to crops: None.

Narrative: The A. thaliana experiments were concluded today. A full summary of the experiments is available in mission summary. Two particular simulants stood as the most successful. Approximately 500 mL of leftover Hoagland’s #2 hydroponic nutrient solution was used on the mystery greens to help in their seedling phase as well as on some of the tomato plants.

The door to the GreenHab was opened to allow for ambient airflow at 1110 hours and then closed at 1200 due to a temperature drop. The door was opened again at 1500 hours because the temperature had spiked due to the intensified sunlight reflecting off of the snow. The door was closed again at 1600 hours for the remainder of the day.

Harvest: None

Support/supplies needed: None.

Copyright © The Mars Society. All rights reserved. | Main Site